One plane at 520 feet is ascending at a rate of 40 feet per minute, while another plane at 3800 feet is descending at a rate of 120 feet per minute. How long will it take the two planes to be at the same altitude?
Answer by ankor@dixie-net.com(22740)

One plane at 520 feet is ascending at a rate of 40 feet per minute, while another plane at 3800 feet is descending at a rate of 120 feet per minute.
How long will it take the two planes to be at the same altitude?
:
Let = time (in minutes) for this to be true
;
40t + 520 = 3800 - 120t
40t + 120t = 3800 - 520
160t = 3280
t = 3280/160
t = 20.5 minutes
:
:
Check this, find at what altitude they are equal
40(20.5) + 520 = 3800 - 120(20.5)
820 + 520 = 3800 - 2460
1340 = 1340